# Deep Sub-Ensembles Meets Quantile Regression: Uncertainty-aware Imputation for Time Series

### v3

@@ 4.3 Bidirectional Recurrent Imputation with Quantile Sub-Ensembles
In this section, we introduce a Bidirectional Long Short-Term Memory network (BiLSTM) enhanced with Quantile Sub-Ensembles for time series imputation. The BiLSTM processes input data in both forward and backward directions simultaneously, allowing it to capture contextual information from both past and future data points. This bidirectional approach makes the model well-suited for time series imputation. The overall architecture of the model is detailed in Figure 3.
Figure 3: The overview of the model framework. The multivariate time series $X$ with missing values are fed to the complement layer which fills missing values. Next, the output of the complement layer $x_{t}^{co}$ and the previous hidden state $h_{t-1}^{f},h_{t+1}^{b}$ is delivered to the feature-based estimation layer to capture the feature relationship. Then we combine $\hat{z}_{t}$ and the temporal decay factor $\gamma_{t}$ which gradually diminishes history information over time as the input to $N$ task networks corresponding to $N$ ensemble members. In the meanwhile, quantile regression loss $L$ will be computed to update the parameters through the process of backpropagation. Finally, the imputed time series of each ensemble is combined and fed into BiLSTM with cell state $c_{t-1}^{f},c_{t+1}^{b}$.
In our case, on account of missing values in time series $X$, we can’t denote $X_{t}$ directly as the input to the model. The solution to this problem is to use a "complement" input. In this initial step, we don’t need to focus too much on the accuracy of the imputed values. It’s sufficient to use a linear layer:
$\hat{x}_{t}$ $=$ $W_{x}h_{t-1}+b_{x},$ (9)
$x_{t}^{co}$ $=$ $m_{t}\odot x_{t}+(1-m_{t})\odot\hat{x}_{t},$ (10)
where $x_{t}^{co}$ is the "complement" time series data and it’s dimension is $n\times d$ ($n$ is the size of batch and d is the dimension of features).
It has been observed that when there is a long time gap between the last observation and the current one, the missing values tend to be replaced by default values. Additionally, if variables are missing for a substantial period of time, the impact of the input variables gradually diminishes over time. Therefore, we introduce a temporal decay factor $\gamma_{t}$ [11] to consider the following important factors:
$\gamma_{t}=\exp\{-max(0,W_{\gamma}\delta_{t})+b_{\gamma}\},$ (11)
@@
In order to utilize feature dimension information during imputation, we introduce feature-based estimation $\hat{z}_{t}$  [10]:
@@
Based on it, we can combine the historical-based estimation $\hat{x}_{t}$ and the feature-based estimation $\hat{z}_{t}$. Then we define $\hat{v}_{t,1},\hat{v}_{t,2},...,\hat{v}_{t,N}$ as the vectors of task networks:
@@
$\hat{v}_{t,i}=\beta_{t,i}\odot\hat{z}_{t}+(1-\beta_{t,i})\odot\hat{x}_{t},$ (14)
where $\beta_{t,1},\beta_{t,2},...,\beta_{t,N}$ is the weights. The function of the task networks layers (12) and (13) is to combine the historical-based estimation $\hat{x}_{t}$ and the feature-based estimation $\hat{z}_{t}$. Adding some layers between (12) and (13) will not promote the information combination, but increase the computational overhead. We can replace missing values in $x_{t}$ with the corresponding values in $\hat{v}_{t,i}$:
@@
$L(q_{i},x_{t},m_{t},v_{t,i})=\sum_{i=1}^{N}\sum_{t=1}^{T}(q_{i}*|\max(x_{t}-v_{t,i},0)|+(1-q_{i})*|\max(0,v_{t,i}-x_{t})|)\odot m_{t}.$ (16)
Then we feed $\overline{v_{t}}$, the mean of $v_{t,1},v_{t,2},...,v_{t,N}$, to the next bidirectional recurrent long short-term memory networks(BiLSTM) to learn the mixed information of all features and capture the short and long relationships of different variables. BiLSTM networks include both a forward and a backward LSTM unit. Given an input sequence $\{\overline{v_{1}},\overline{v_{2}},...,\overline{v_{T}}\}$, the computations are:
  • Forward LSTM:
    $h_{t}^{f}=\text{LSTM}_{f}(\overline{v_{t}},h_{t-1}^{f},c_{t-1}^{f})$
  • Backward LSTM:
    $h_{t}^{b}=\text{LSTM}_{b}(\overline{v_{t}},h_{t+1}^{b},c_{t+1}^{b})$
Here, $\text{LSTM}_{f}$ and $\text{LSTM}_{b}$ denote the forward and backward LSTM units, $h_{t}^{f}$ and $h_{t}^{b}$ are the hidden states of the forward and backward, $c_{t}^{f}$ and $c_{t}^{b}$ are the cell states. The final Bi-LSTM output is typically the concatenation of the forward and backward hidden states:
$h_{t}=[h_{t}^{f};h_{t}^{b}]$ (17)
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] represents the concatenation operation. The overall procedure of our algorithm is summarized in Algorithm 1.
@@
Multivariate time series $X=\{x_{1:T}^{1:K}\}\in\mathbb{R}^{T\times K}$, number of ensmeble members $N$, quantiles $\{q_{1},q_{2},...,q_{N}\}$, iterations $T$.
@@
     $\hat{x}_{t}=W_{x}h_{t-1}+b_{x}$
6:     $x_{t}^{co}=m_{t}\odot x_{t}+(1-m_{t})\odot\hat{x}_{t}$triangleright fill in the missing values beforehand
     computing mean absolute error loss: $L_{1}(x_{t}^{co},x_{t},m_{t})$
8:     introduce a temporal decay factor $\gamma_{t}=exp\{-max(0,W_{\gamma}\delta_{t})+b_{\gamma}\}$
@@
10:     computing mean absolute error loss: $L_{2}(\hat{z}_{t},x_{t},m_{t})$
@@
         define combined vectors $\hat{v}_{t,i}=\beta_{t,i}\odot\hat{z}_{t}+(1-\beta_{t,i})\odot\hat{x}_{t}$
14:         $v_{t,i}=m_{t}\odot x_{t}+(1-m_{t})\odot\hat{v}_{t,i}$triangleright replace missing values
@@
16:     $L(q_{i},x_{t},m_{t},v_{t,i})=\sum_{1}^{N}\sum_{1}^{T}(q_{i}*|max(x_{t}-v_{t,i},0)|+(1-q_{i})*|max(0,v_{t,i}-x_{t})|)\odot m_{t}$
     minimize $L_{1}(x_{t}^{co},x_{t},m_{t})+L_{2}(\hat{z}_{t},x_{t},m_{t})+L(q_{i},x_{t},m_{t},v_{t,i})$
18:     compute $\overline{v_{t}}$, the mean of $v_{t,1},v_{t,2},...,v_{t,N}$ , $\overline{v_{t}}$ is the imputed value
     $h_{t}^{f}=\text{LSTM}_{f}(\overline{v_{t}},h_{t-1}^{f},c_{t-1}^{f}),h_{t}^{b}=\text{LSTM}_{b}(\overline{v_{t}},h_{t+1}^{b},c_{t+1}^{b})$
20:     $h_{t}=[h_{t}^{f};h_{t}^{b}]$triangleright LSTMs
end for
4.4 Theoretical Discussions
Limitation of quantiles. The Bayesian Gaussian Mixture model employed for calculating quantiles utilizes a maximum likelihood estimation on the posterior distribution, thereby achieving convergence to the desired $N$ quantiles. However, increasing the number of quantiles corresponds to a greater overlap between the intensity distributions of adjacent quantile intervals which can potentially lead to overfitting. This occurs because the model becomes more complex with more quantiles, potentially fitting noise in the data rather than capturing the underlying relationship [8]. As the number of quantiles increases, the model’s flexibility grows, which can lead to higher variance and poorer generalization.
Robustness with high missing rates. When dealing time series data with a high missing rate, RNN-based networks can capture temporal dependencies from the non-missing parts of the time series to obtain periodic features. Since real-world time series often exhibit significant periodicity, we can leverage this periodicity to impute large consecutive missing segments accurately [48]. Meanwhile, the proposed Quantile Sub-Ensembles method alleviate overfitting to the mean sample intensity and extreme values by fitting each quantile-regression-based task network to the corresponding quantile interval, which provide robust and reliable imputations.
@@ 5 Experiments
In this section, we compare the proposed method with existing methods in terms of imputation accuracy and uncertainty on time series. We conduct experiments on two public datasets: heath-care and air quality from relevant application domains. Subsequently, we also show the hyper-parameters sensitivity and computation efficiency of the proposed method. These experimental settings, including datasets, hyperparameters and models, are compatible with the most of the baselines.
@@ 5.1 Dataset Description
Heath-care data: Heath-care data in PhysioNet Challenge 2012 [41] consists of 4000 clinical time series with 35 variables. Each time series is recorded within the first 48 hours of the patient’s admission to the ICU, so we process it hourly with 48 time steps. Since the data is irregularly sampled, the missing rate is up to $80\%$ in total. For this dataset, we randomly eliminate 10/50/90% of observed values and regard them as the ground-truth, which is only used for testing the imputation.
Air quality data: Air quality data contains the PM2.5 readings from 36 air quality monitoring stations in Beijing from 2014/05/01 to 2015/04/30. Overall, the missing rate is 13.3%. Following previous studies [10; 42], we set 36 consecutive time steps as one-time series and use $3^{rd}$, $6^{th}$, $9^{th}$ and $12^{th}$ months data as a test set and the rest for training. The dataset’s missing patterns are resolved thanks to the inclusion of artificial ground-truth (The air quality dataset includes artificially recorded ground-truth values and this specific phrase is adopted from [45]). Therefore, we have no need to choose different proportions of the data to train and test the model.
@@ 5.2 Experiment Settings
Following the previous setting [10], we use an Adam optimizer with a learning rate of 0.001 and batch size 32 to train our model, which follows previous studies. For all datasets, we normalize the data of all features at time dimension to acquire zero mean and unit variance, which can make model training stable. In addition, we set 5 quantile levels $Q_{1}=[0.1,0.25,0.5,0.75,0.9]$ to calculate the quantile regression loss of each ensemble in parallel.
@@ 5.3 Baselines
$\bullet$ Multitask GP [7]: Multi-task Gaussian processes (Multitask GP) learn the covariance between time points and features simultaneously, boosting imputation performance and allowing for flexible modeling of task dependencies.
$\bullet$ GP-VAE  [20]: GP-VAE assumes that a high-dimensional time series can be represented by a low-dimensional Gaussian process that evolves smoothly over time. It reduces non-linear dimensionality and handles missing data with VAE.
$\bullet$ RDIS [14]: RDIS is a novel training method for time series imputation. RDIS applies a random drop on the missing data to learn more by imputing the random drop values.
$\bullet$ V-RIN [34]: V-RIN is a variational-recurrent imputation network that combines an imputation and a prediction network by considering the relevant characteristics, temporal dynamics and uncertainty.
$\bullet$ BRITS [10]: BRITS adopts a bidirectional recurrent network to capture temporal dependencies of time series for imputing missing values. Moreover, in order to make full use of the relationship of features, It combines the historical-based estimation with the feature-based estimation. Finally, a classification task is designed to confirm the imputed values.
$\bullet$ CSDI [45]: CSDI employs score-based diffusion models conditioned on observations. Its probability calculation process can effectively quantify uncertainty and make probabilistic imputation for time series data.
$\bullet$ CSBI [12]: CSBI addresses the problem of time series imputation using the Schrödinger bridge problem (SBP), which is gaining popularity in generative modeling. It provides a convergence analysis for the Schrödinger bridge algorithm based on approximated projections, bridging a theoretical gap. Practically, SBP is applied to probabilistic time series imputation, generating missing values conditioned on observed data.

### v4

@@ 3.3 Bidirectional Recurrent Imputation with Quantile Sub-Ensembles
In this section, we introduce a Bidirectional Long Short-Term Memory network (BiLSTM) enhanced with Quantile Sub-Ensembles for time series imputation. The BiLSTM processes input data in both forward and backward directions simultaneously, allowing it to capture contextual information from both past and future data points. By incorporating Quantile Sub-Ensembles, the model estimates multiple quantiles that capture uncertainty in the data, leading to more reliable imputation performance and reducing the risk of overfitting in complex time series scenarios. The overall architecture of the model is detailed in Fig. 3.
We denote the input multivariate time series as $X=\{x_{1:T}^{1:K}\}\in\mathbb{R}^{T\times K}$ where $T$ is the length of the time series, $K$ is the feature dimension, and $x_{t}^{k}$ represents the $k$-th feature value at the timestamp $d_{t}$. To represent the missing values in $X$, we define a mask matrix $M=\{m_{1:T}^{1:K}\}\in\mathbb{R}^{T\times K}$, where $m_{t}^{k}=1$ if $x_{t}^{k}$ is observed and $m_{t}^{k}=0$ if $x_{t}^{k}$ is missing. In the time series imputation task, our target is to fit an imputation function $F$: $X\longrightarrow\hat{X}$, where $\hat{X}=\{\hat{x}_{1:T}^{1:K}\}\in\mathbb{R}^{T\times K}$ is the output imputed time series.
$\delta_{t}^{k}=\begin{cases}0,&\text{if }t=1,\\ d_{t}-d_{t-1}+\delta_{t-1}^{k},&\text{if }t>1\text{ and }m_{t-1}^{k}=0,\\ d_{t}-d_{t-1},&\text{if }t>1\text{ and }m_{t-1}^{k}=1.\end{cases}$
In this paper, we focus on probabilistic time series imputation [GPVAE], i.e., estimating the probabilistic distribution of time series missing values through observation. Therefore, uncertainty can be quantified while imputing missing values.
Figure 3: The overview of the model framework. The multivariate time series $X$ with missing values are fed to the complement layer which fills missing values. Next, the output of the complement layer $x_{t}^{co}$ and the previous hidden state $h_{t-1}^{f},h_{t+1}^{b}$ is delivered to the feature-based estimation layer to capture the feature relationship. Then we combine $\hat{z}_{t}$ and the temporal decay factor $\gamma_{t}$ which gradually diminishes history information over time as the input to NN task networks corresponding to NN ensemble members. In the meanwhile, quantile regression loss $L$ is computed to update the parameters through the process of backpropagation. Finally, the imputed time series of each ensemble is combined and fed into BiLSTM with cell state $c_{t-1}^{f},c_{t+1}^{b}$.
In our case, on account of missing values in time series $X$, we can’t denote $X_{t}$ directly as the input to the model. The solution to this problem is to use a “complement” input. In this initial step, we don’t need to focus too much on the accuracy of the imputed values. It’s sufficient to use a linear layer:
$x_{t}^{\prime}$ $=$ $W_{x}h_{t-1}+b_{x},$ (9)
$x_{t}^{co}$ $=$ $m_{t}\odot x_{t}+(1-m_{t})\odot x_{t}^{\prime},$ (10)
where $x_{t}^{co}$ is the “complement” time series data and it’s dimension is $n\times d$ ($n$ is the size of batch and d is the dimension of features).
It has been observed that if variables are missing for a substantial period of time, the influence of the input variables will fade away over time. Therefore, we introduce a temporal decay factor $\gamma_{t}$ [2018Recurrent] to consider the following important factors:
$\gamma_{t}=\exp\{-\max(0,W_{\gamma}\delta_{t})+b_{\gamma}\},$ (11)
@@
In order to utilize feature dimension information during imputation, we introduce feature-based estimation $\hat{z}_{t}$  [BRITS]:
@@
Based on it, we can combine the historical-based estimation $x_{t}^{\prime}$ and the feature-based estimation $\hat{z}_{t}$. Then we define $\hat{v}_{t,1},\hat{v}_{t,2},...,\hat{v}_{t,N}$ as the vectors of task networks:
@@
$\hat{v}_{t,i}=\beta_{t,i}\odot\hat{z}_{t}+(1-\beta_{t,i})\odot x_{t}^{\prime},$ (14)
where $\beta_{t,1},\beta_{t,2},...,\beta_{t,N}$ is the weights. The function of the task networks layers (12) and (13) is to combine the historical-based estimation $x_{t}^{\prime}$ and the feature-based estimation $\hat{z}_{t}$. Adding some layers between (12) and (13) will not promote the information combination, but increase the computational overhead. We can replace missing values in $x_{t}$ with the corresponding values in $\hat{v}_{t,i}$:
@@
[…]
Then we feed $\overline{v_{t}}$, the mean of $v_{t,1},v_{t,2},...,v_{t,N}$, to the next bidirectional recurrent long short-term memory networks(BiLSTM) to learn the mixed information of all features and capture the short and long relationships of different variables. The overall procedure of our algorithm is summarized in Algorithm 1.
3.4 Theoretical Discussions
Limitation of quantiles. The Bayesian Gaussian Mixture model employed for calculating quantiles utilizes a maximum likelihood estimation on the posterior distribution, thereby achieving convergence to the desired $N$ quantiles. However, increasing the number of quantiles corresponds to a greater overlap between the intensity distributions of adjacent quantile intervals which can potentially lead to overfitting. This occurs because the model becomes more complex with more quantiles, potentially fitting noise in the data rather than capturing the underlying relationship [fakoor2023flexible]. As the number of quantiles increases, the model’s flexibility grows, which can lead to higher variance and poorer generalization.
@@ 3.4 Theoretical Discussions
Multivariate time series $X=\{x_{1:T}^{1:K}\}\in\mathbb{R}^{T\times K}$, number of ensmeble members NN, quantiles $\{q_{1},q_{2},...,q_{N}\}$, iterations $T$.
@@
  $x_{t}^{\prime}=W_{x}h_{t-1}+b_{x}$
6:  $x_{t}^{co}=m_{t}\odot x_{t}+(1-m_{t})\odot x_{t}^{\prime}$ $\triangleright$ Fill in the missing values beforehand
  Compute mean absolute error loss: $L_{1}(x_{t}^{co},x_{t},m_{t})$
8:  Introduce a temporal decay factor $\gamma_{t}=\exp\{-\max(0,W_{\gamma}\delta_{t})+b_{\gamma}\}$
@@
10:  Compute mean absolute error loss: $L_{2}(\hat{z}_{t},x_{t},m_{t})$
@@
   Define combined vectors $\hat{v}_{t,i}=\beta_{t,i}\odot\hat{z}_{t}+(1-\beta_{t,i})\odot x_{t}^{\prime}$
14:   $v_{t,i}=m_{t}\odot x_{t}+(1-m_{t})\odot\hat{v}_{t,i}$ $\triangleright$ Replace missing values
@@
16:  $L(q_{i},x_{t},m_{t},v_{t,i})=\sum_{1}^{N}\sum_{1}^{T}(q_{i}*|\max(x_{t}-v_{t,i},0)|+(1-q_{i})*|\max(0,v_{t,i}$-
  $x_{t})|)\odot m_{t}$
18:  Minimize $L_{1}(x_{t}^{co},x_{t},m_{t})+L_{2}(\hat{z}_{t},x_{t},m_{t})+L(q_{i},x_{t},m_{t},v_{t,i})$
  Compute $\overline{v_{t}}$, the mean of $v_{t,1},v_{t,2},...,v_{t,N}$ , $\hat{x}_{t}=\overline{v_{t}}$ is the imputed value
20:  $h_{t}^{f}=\text{LSTM}_{f}(\overline{v_{t}},h_{t-1}^{f},c_{t-1}^{f}),h_{t}^{b}=\text{LSTM}_{b}(\overline{v_{t}},h_{t+1}^{b},c_{t+1}^{b})$
  $h_{t}=[h_{t}^{f};h_{t}^{b}]$ $\triangleright$ LSTMs
22:end for
Robustness with high missing rates. When dealing with time series data with a high missing rate, RNN-based networks can capture temporal dependencies from the non-missing parts of the time series to obtain periodic features. Since real-world time series often exhibit significant periodicity, we can leverage this periodicity to impute large consecutive missing segments accurately [cyclernn]. Meanwhile, the proposed Quantile Sub-Ensembles method alleviate overfitting on the mean sample intensity and extreme values by fitting each quantile-regression-based task network to the corresponding quantile interval, which provide robust and reliable imputations.
@@ 4 Experiments
In this section, we compare the proposed method with existing methods in terms of imputation accuracy and uncertainty on time series. We conduct experiments on five public datasets (solar, traffic, healthcare, electricity, and air-quality) from relevant application domains. Subsequently, we also show the hyper-parameters sensitivity and computation efficiency of the proposed method.
@@ 4.1 Dataset Description
Solar dataset [^1] https://www.nrel.gov/grid/solar-power-data.html contains power production measurements from 137 photovoltaic (PV) plants located in Alabama, totaling 52,560 timestamps. Traffic dataset [traffic] comprises the traffic speed of 214 features over a five-day period in Guangzhou, with a sampling rate of one reading per 10 minutes. Electricity dataset [^2] https://archive.ics.uci.edu/dataset/321/electricityloaddiagrams20112014 records the hourly electricity consumption (kWh) of 370 clients from January 2011 to December 2014. Air-quality dataset [^3] https://archive.ics.uci.edu/dataset/501/beijing+multi+site+air+quality+data[air-quality] comprises hourly air pollutant observations from 12 monitoring sites in Beijing, spanning March 2013 to February 2017. For each site, 11 continuous time series variables are recorded. Healthcare dataset from PhysioNet Challenge 2012 [2013Predicting] consists of 4000 clinical time series with 35 variables. Each time series is recorded within the first 48 hours of the patient’s admission to the ICU, so we process it hourly with 48 time steps. Due to the incomplete data collection, solar, air-quality, and healthcare datasets are inherently missing. For example, missing records are often caused by sensor malfunctions. This is likely closer to the MAR (Missing at Random) pattern. The traffic and electricity datasets is complete. For five datasets, we select different missing rates of the available data points as observations under the MCAR (Missing Completely at Random) pattern for model training, and the rest to serve as the ground truth for testing.
@@ 4.2 Experiment Settings
Following the previous setting [CSDI], we use an Adam optimizer with a learning rate of 0.001 and batch size 32 to train our model, which follows previous studies. For all datasets, we normalize the data of all features at time dimension to acquire zero mean and unit variance, which can make model training stable. In addition, we set 5 quantile levels $Q_{1}=[0.1,0.25,0.5,0.75,0.9]$ to calculate the quantile regression loss of each ensemble in parallel.
@@ 4.3 Baselines
  • $\bullet$
    SAITS [SAITS] employs a diagonally masked self-attention mechanism to capture temporal dependencies and adopts a joint optimization approach during training.
  • $\bullet$
    TimesNet [timesnet] addresses the limited representational capacity of 1D time series by transforming them into 2D tensors across multiple periods, thereby extending temporal variation analysis into the 2D space.
  • $\bullet$
    Multitask GP [2008Multi] learns the covariance between time points and features simultaneously, boosting imputation performance and allowing for flexible modeling of task dependencies.
  • $\bullet$
    GP-VAE  [GPVAE] assumes that a high-dimensional time series can be represented by a low-dimensional Gaussian process that evolves smoothly over time. It reduces non-linear dimensionality and handles missing data with VAE.
  • $\bullet$
    V-RIN [V-RIN] is a variational-recurrent imputation network that combines an imputation and a prediction network by considering the relevant characteristics, temporal dynamics and uncertainty.
  • $\bullet$
    CSDI [CSDI] employs score-based diffusion models conditioned on observations. Its probability calculation process can effectively quantify uncertainty and make probabilistic imputation for time series data.
  • $\bullet$
    CSBI [CSBI] addresses the problem of time series imputation using the Schrödinger bridge problem (SBP), which is gaining popularity in generative modeling.
In addition, we compare the proposed method with the foundational Forward Filling (Forward), Linear Interpolation (Linear), and BiLSTM to demonstrate the superior effectiveness of our method.
Table 1: Performance comparison of MAE and CRPS at 50% missing rate.
solar traffic healthcare electricity air-quality
Model MAE CRPS MAE CRPS MAE CRPS MAE CRPS MAE CRPS
Forward 2.212 - 7.791 - 0.923 - 1.032 - 0.994 -
Linear 1.975 - 5.817 - 0.453 - 0.734 - 0.245 -
SAITS 1.327 - 3.291 - 0.341 - 0.876 - 0.194 -
TimesNet 1.221 - 3.205 - 0.414 - 0.305 - 0.210 -
BiLSTM 0.845 - 3.197 - 0.348 - 0.329 - 0.169 -
Multitask GP 1.706 0.203 3.532 0.092 0.516 0.581 1.066 0.913 0.313 0.342
GP-VAE 1.811 0.368 3.419 0.084 0.478 0.774 1.097 0.862 0.258 0.294
V-RIN 1.547 0.362 3.375 0.107 0.365 0.831 0.541 0.583 0.197 0.268
CSDI 1.104 0.166 3.202 0.076 0.315 0.338 0.278 0.247 0.156 0.213
CSBI 1.033 0.153 3.182 0.074 0.318 0.352 0.263 0.225 0.172 0.236
Proposed method 0.568 0.085 2.884 0.067 0.313 0.378 0.212 0.157 0.154 0.191